Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Wednesday, November 11, 1998

The following is a brief summary of the environmental monitoring activities that were completed on November 11, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

11/11/98 (Wednesday)

Segment 1 (Emidio to ANF). Report for 11/11/98. Rain fell throughout most of the workday. Most operations were shut down mid-day. Tie-in trench was excavated along Lebec Road at EMP 25.21. Clearing operations continued along the skipped area at EMP 21.73 to EMP 21.52. Tie-in work near EMP 18.48 continued until shut down, due to rain. Most other work on Tejon Bench stopped by 1430 hrs. Excavation along Grapevine Road continued northward to ~ EMP 15.97, where the trench began to fill with water from run-off from the sloping terrain. Backfill and Compaction along Aqueduct Access Road B continued to ~ EMP 6.5. Road work and repairs around manholes were made at several locations along Access Road A. On Hwy 166, Excavation and trench shoring continued. Pipe was lowered-in at ~ EMP .03. The trench began to fill with water in the afternoon. Station work was concentrated at Grapevine station, with electrical and welding operations. No crews were present at Emidio station during my monitoring there. Work continued at The California Aqueduct bore site, the Tecuya Creek bore and I-5 Bore sites. Pipe Bending operations took place at the top of Grapevine Road, EMP 16.20.

Segment 2 (ANF to Sylmar) Report for 11/11/98: 2C (Templin Hwy to Castaic): Overnight rain made the ROW difficult for travel in Marple Canyon, the location for most current construction activity on this Segment. Work today was limited to a few revegetation activities. Rains beginning mid-day ended that work for the day. 2B (Castaic to Sylmar): 240 feet of boring was completed at the Feedmill Rd directional drill. At Calgrove, backfilling of bore pits continued. Also, already excavated trench was cleaned out. At Sierra Hwy 150 feet of trench was excavated into CalTrans area.. South of Sierra Hwy, pipe was lowered-in, welded, x-rayed, and coated. At Sunshine Cyn, base rock was placed around existing pipe, slurry poured, and roadway repaired.

Segment 3A, 3B, & 3C (Sylmar to Los Angeles): Report for 11 November 1998: 3A: Backfill, padding, and fiber optic conduit crews focused on the north side of the San Fernando Bridge crossing, MP 86.76. Welding, x-ray, and coating occurred at the tie in to the Hollywood Way Bridge crossing, MP 85.60. Work on the Tuxford bore site continued. Pipeline markers continued to be installed in various areas along the alignment. 3B: No work occurred. 3C: The bore at Hubbard Avenue was completed. The crew will move to the Tuxford bore. Welding, x-ray, and coating worked on the pipe in San Fernando Road, near the I-5 overpass. Crews worked late to slurry the pipe and fiber optic conduits because the trench was collapsing, undermining the roadway. Welding and dye checks were completed on the pipe over the aqueduct.

Segment 4B(El Segundo to Lynwood): Due to the holiday (Veterans Day) only the crew at Western Bore pit ESEMP 5.56 was working today. This crew was excavating to remove the damage pipe.

Segment 5 (Lynwood to Long Beach) Report for 11 November 1998: The horizontal directional drill at Highway 91/Compton Creek resumed without problem. The frac-outs of yesterday were contained and removed. The first ream was almost completed today. The bore crew at Wilmington are progressing without problems.

Inspection Summary

In the past 24 hours, 5 inspection event(s) were reported for 5 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page